# JanusVLN: Decoupling Semantics and Spatiality with Dual Implicit Memory for Vision-Language Navigation

### v1

@@ Appendix B More ablation studies
Appendix C More qualitative results
This section presents further qualitative analysis of JanusVLN in both real-world and simulated environments. For real-world settings in Figure 5, we selected navigation tasks that involve simple and complex instructions, diverse sites, and spatial understanding, where JanusVLN demonstrates excellent generalization. For simulated environments in Figure 6 and 7, we chose complex trajectories and long instructions from the unseen validation sets of R2R-CE and RxR-CE. Leveraging its dual implicit memory, JanusVLN effectively follows these instructions to complete challenging navigation tasks.

### v2

@@ Appendix C More ablation studies
Appendix D Statistical analysis
Figure 7: Performance on spatial understanding tasks.
Success and strengths analysis.
In Figure 7, We measured the success rate on instructions requiring spatial understanding (i.e., those containing terms like ’farthest,’ ’nearest,’ ’larger,’ ’smaller,’ ’rightmost,’ ’leftmost,’ ’first,’ ’second,’ ’front,’ ’back,’ etc.). We find that the superiority of JanusVLN over prior methods is more pronounced in scenarios requiring spatial understanding than its average gain across all tasks, demonstrating its strong spatial awareness.
Figure 8: Performance on various instruction lengths/complexity.
Performance by instruction length.
We analyzed the trends in SR and SPL for both StreamVLN and JanusVLN as instruction length increases in Figure 8. Both models achieve high SR and SPL on relatively simple instructions (1-150 words). However, their performance declines on moderately complex instructions (150-400 words), indicating a need to enhance the models’ ability to decompose and comprehend complex directives. For the most complex instructions (400-550 words), StreamVLN’s performance continues to degrade, eventually reaching zero. In contrast, JanusVLN’s performance improves, benefiting from its dual implicit memory paradigm. This is likely because these lengthy instructions provide highly detailed, step-by-step guidance that the model can effectively follow.
Figure 9: Visualization and presentation of the types of failure cases.
Failure case analysis.
Our statistical analysis reveals two predominant types of failure cases for JansuVLN in Figure 9. First, when the agent deviates from the optimal trajectory, it attempts to correct its course but often fails to recover, leading to compounding errors and eventual failure. Although we collected a limited amount of non-optimal trajectory data via DAgger, it is insufficient to enable robust error correction. Second, JanusVLN appears to employ an overly aggressive stopping policy, sometimes halting prematurely upon sighting the destination and thus failing to enter the success radius. This may be because the spatial information from its VGGT encoder lacks real-world scale, resulting in inaccurate distance estimation.
Appendix E More qualitative results
Visualization analysis of spatial geometric tokens.
We demonstrate how spatial geometry tokens aid navigation by visualizing them as depth maps and point clouds in Figure 10. In the first example, the depth map derived from the tokens captures precise depth information, enabling a more accurate localization of the farthest chair. In the second, the point cloud constructed from the tokens clearly reveals the chair behind the sink counter. In the third example, both visualizations distinctly represent the size of the door. Finally, in the fourth example, visualizations reveal that the tokens focus on the rightmost house, as reflected in both its depth map and point cloud. In conclusion, the spatial information captured by these tokens is crucial for spatial understanding.
More qualitative results.
This section presents further qualitative analysis of JanusVLN in both real-world and simulated environments. For real-world settings in Figure 11, we selected navigation tasks that involve simple and complex instructions, diverse sites, and spatial understanding, where JanusVLN demonstrates excellent generalization. For simulated environments in Figure 12 and 13, we chose complex trajectories and long instructions from the unseen validation sets of R2R-CE and RxR-CE. Leveraging its dual implicit memory, JanusVLN effectively follows these instructions to complete challenging navigation tasks.
Figure 10: The analysis on the effectiveness of spatial gemetric tokens.